# Generate-on-Graph: Treat LLM as both Agent and KG in Incomplete Knowledge Graph Question Answering

### v1

@@ Abstract
To address the issue of insufficient knowledge and the tendency to generate hallucination in Large Language Models (LLMs), numerous studies have endeavored to integrate LLMs with Knowledge Graphs (KGs). However, all these methods are evaluated on conventional Knowledge Graph Question Answering (KGQA) with complete KGs, where the factual triples involved in each question are entirely covered by the given KG. In this situation, LLM mainly acts as an agent to find answer entities by exploring the KG, rather than effectively integrating internal and external knowledge sources. However, in real-world scenarios, KGs are often incomplete to cover all the knowledge required to answer questions. To simulate real-world scenarios and evaluate the ability of LLMs to integrate internal and external knowledge, in this paper, we propose leveraging LLMs for QA under Incomplete Knowledge Graph (IKGQA), where the given KG doesn’t include all the factual triples involved in each question. To handle IKGQA, we propose a training-free method called Generate-on-Graph (GoG) that can generate new factual triples while exploring on KGs. Specifically, we propose a selecting-generating-answering framework, which not only treat the LLM as an agent to explore on KGs, but also treat it as a KG to generate new facts based on the explored subgraph and its inherent knowledge. Experimental results on two datasets demonstrate that our GoG can solve IKGQA to a certain extent, while almost all previous methods cannot perform well on IKGQA.
@@ 1 Introduction
Figure 1: An example of Incomplete Knowledge Graph Question Answering (IKGQA), where yellow and red nodes represent topic and answer entity, respectively. Red dash line represents this triple (Cupertino, timezone, Pacific Standard Time) is missing in the KG.
@@
Large Language Models (LLMs) Brown et al. (2020); Bang et al. (2023) have made great success in various natural language processing (NLP) tasks. Benefiting from extensive model parameters and vast amounts of pre-training corpus, LLMs can solve complex reasoning tasks through In-Context Learning (ICL) Dong et al. (2023), without fine-tuning for specific tasks.
@@
However, LLMs are still suffer from insufficient knowledge and the tendency to generate hallucinationHuang et al. (2023); Li et al. (2023a). To mitigate this issue, many methods that incorporate LLMs with Knowledge Graphs (KGs) Ji et al. (2021) has been proposed Pan et al. (2023), where KGs provide accurate and abundant factual knowledge in triple format while LLMs provide strong natural language processing ability. These works can be roughly divided into two categories: (1) Semantic Parsing (SP) methods Li et al. (2023c); Nie et al. (2023), which use LLMs to generate logical forms with ICL, and then obtain answers by executing these logical queries on KGs. (2) Retrieval Augmented (RA) methods Li et al. (2023d), which retrieve information related to the question from KGs as external knowledge to help LLMs to obtain the answers. The retrieval format can be triples Li et al. (2023d), paths Sun et al. (2023); Luo et al. (2023b) and subgraphs Wang et al. (2023a).
Semantic parsing based methods exclusively treat LLMs as parser, not only do they ignore LLMs’ inherent knowledge and reasoning ability, but they also depend heavily on KGs’ quality and completeness Sun et al. (2023). Therefore, more attention is paid to retrieval augmented methods.
Figure 2: The performance of ToG Sun et al. (2023) and our GoG under complete/incomplete KG. Dash line indicates the performance of LLM with CoT Wei et al. (2023) prompting (without KG).
@@
Although these retrieval augmented methods claim to solve the drawbacks of semantic parsing methods and obtain good performance on conventional Knowledge Graph Question Answering (KGQA) Yih et al. (2016a), it is still hard to determine if they effectively integrate knowledge from KGs and LLMs, and how much reasoning ability is utilized. One crucial reason is that, in conventional KGQA tasks, the factual triplets involved in each question are entirely covered by the KG. Therefore, in this scenario, LLMs still play the role of parser which only needs to identify the relationship path starting from the topic entity to the answer entity. For example, for the question "What is the timezone of the area where Apple headquarters is located?" in Figure 1, the LLM only needs to start from Apple headquarters, sequentially choose the relations located_in and timezone to find the answer. That means LLMs only need to ground the relationship appearing in the question to the specific relation in the KG to reach the answer entity Pacific Standard Time. Therefore, in conventional KGQA, LLM mainly acts as an agent to find answer entities by exploring the KG, rather than effectively integrating internal and external knowledge sources.
@@
However, in real-world scenarios, KGs are often incomplete to cover all the knowledge required to answer questions. Besides, LLMs contain rich knowledge content and have powerful reasoning ability. Therefore, to evaluate the ability of LLMs to integrate internal and external knowledge, in this paper, we propose leveraging LLMs for QA under incomplete KG (IKGQA). The distinction between IKGQA and conventional KGQA is that IKGQA does not encompass all the factual triplets relevant to each question, rendering the KG in IKGQA incomplete. For example, as shown in Figure 1, for answering the same question "What is the timezone of the area where Apple headquarters is located?", the key triple (Cupertino, timezone, Pacific Standard Time) does not exist in the KG. This means that even if the correct SPARQL query is generated, it may not retrieve the final answer  [^1] SP methods do not parse ”timezone” into two hop path ”located_in -¿ timezone” because in the training set, ”timezone” only corresponds to one hop path ”timezone”, more details can be found in Appendix A.. Compared to KGQA, IKGQA holds greater research significance for the following reasons: (1) it is closer to real-world scenarios where the given KG is incomplete to answer users’ questions. (2) it evaluates LLMs’ reasoning ability and its capability to integrate inherent and external knowledge.
As shown in Figure 2, although previous methods have achieved outstanding performance in KGQA, their performance has significantly declined in IKGQA, even is worse than that without KG. This discrepancy suggests that these methods might not fully integrate the external and inherent knowledge of LLMs as professed.
To address the challenges in IKGQA, we propose a novel method called Generate-on-graph (GoG), which adopts a selecting-generating-answering framework consisting of three main steps: (1) Selecting, where LLMs select the relations most relevant to the current question and expanding the subgraph using these relations. (2) Generating, where LLMs use its inherent knowledge and reasoning abilities to generate new factual triples based on the explored subgraph. For example, if the KG already contains the triples (Cupertino, located_in, California), (California, timezone, Pacific Standard Time), LLMs can infer new triple (Cupertino, timezone, Pacific Standard Time). (3) Answering, where LLMs try to answer the question based on the retrieval and generated knowledge. If the information is still insufficient, the process is repeated by going back to Steps 1 and 2 until the LLMs can answer the question. The codes and data are available at https://github.com/YaooXu/GoG.
@@
    We propose leveraging LLMs for QA under incomplete KG (IKGQA), which is closer to real-world scenarios and can evaluate LLMs’ ability better, and construct relevant datasets.
@@
    We propose a method called Generate-on-Graph (GoG), which uses the selecting-generating-answering framework, to address IKGQA.
@@
    Experimental results on two datasets show the superiority of GoG, and demonstrate that an incomplete KG can still help LLMs answer complex questions by providing related structured information, even without directly providing the answers.
Figure 3: Comparison of four methods in solving IKGQA: (a) Standard CoT prompting (closed-book, without KG), (b) Semantic parsing based method (e.g., KB-BINDER Li et al. (2023c)), (c) Path retrieval method (e.g., ToG Sun et al. (2023)), (d) The proposed GoG with selecting-generating-answering framework.
@@ 2 Related Work
Question Answering Under Incomplete KG. Some previous works Saxena et al. (2020); Zan et al. (2022) attempt to train KG embeddings to predict answers by similarity scores under incomplete KG. Compared to these previous KGE-based works, we propose leveraging LLMs for QA under incomplete KG to study whether LLMs can integrate internal and external knowledge well.
Unifying KGs and LLMs for KGQA. Various methods have been proposed to unify KGs and LLMs to solve KGQA, these methods can be classified into two categories: Semantic Parsing (SP) methods Li et al. (2023c); Nie et al. (2023) and Retrieval Augmented (RA) methods Luo et al. (2023b); Sun et al. (2023). SP methods transform the question into a structural query using LLMs. These queries can then be executed by a KG engine to derive answers based on KGs Sun et al. (2020). KB-BINDER Li et al. (2023c) generates the drafts as preliminary logical forms first, and then binds the drafts to the executable ones with entity and relation binders. However, the effectiveness of these methods relies heavily on the quality of the generated queries and the completeness of KGs. RA methods retrieve related information from the KG to improve the reasoning performance Li et al. (2023b). ToG Sun et al. (2023) treats the LLM as an agent to interactively explore relation paths step-by-step on KGs and perform reasoning based on the retrieved paths. RoG Luo et al. (2023b) first generates relation paths as faithful plans, and then use them to retrieve valid reasoning paths from the KGs for LLMs to reason. Our GoG belongs to retrieval augmented methods, we also utilize the knowledge modeling ability of LLMs, as well as the semantic parsing ability and reasoning ability.
LLM reasoning with Prompting. Many works have been proposed to elicit the reasoning ability of LLMs to solve complex tasks through prompting Wei et al. (2023); Khot et al. (2023). Complex CoT Fu et al. (2023) creates and refine rationale examples with more reasoning steps to elicit better reasoning in LLMs. Self-Consistency Wang et al. (2023b) fully explores various ways of reasoning to improve their performance on reasoning tasks. DecomP Khot et al. (2023) solves complex tasks by instead decomposing them into simpler sub-tasks and delegating these to sub-task specific LLMs. ReAct Yao et al. (2023) treats LLMs as agents that interact with the environment and make decisions to retrieve information from external source. GoG can be viewed as a fusion of ReAct and DecomP, thereby enabling a more comprehensive utilization of the diverse capabilities inherent in LLMs for addressing complex questions.
@@ 3 Preliminary
In this section, we firstly introduce Knowledge Graph (KGs). Then, we use symbols of KGs to define relation path and Knowledge Graph Question Answering (KGQA).
Knowledge Graphs (KG) is a set of factual triples, i.e., $\mathcal{G}=\{(h,r,t)\in\mathcal{V}\times\mathcal{R}\times\mathcal{V}\}$, where $h,r\in\mathcal{V}$ denote the head and tail entity, $r\in\mathcal{R}$ represents the relation.
Knowledge Graph Question Answering (KGQA) is a reasoning task that aims to predict answer entities $e_{a}\in\mathcal{A}_{q}$ based on $\mathcal{G}$. Following previous work Sun et al. (2019), we call entities mentioned in question $q$ as topic entities, denoted as $e_{t}\in\mathcal{T}_{q}$. Many datasets Talmor and Berant (2018); Yih et al. (2016b) give the standard SPARQL query of each question, which demonstrates a relation path from topic entity $e_{t}$ to answer entity $e_{a}$. We call this path as gold relation path, denote it as $w_{g}=e_{q}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}...\xrightarrow{r_{l}}e_{a}$. For example, the gold relation path of question in Figure 1 is $w_{g}=Apple\;Inc\xrightarrow{headquartered}Cupertino\xrightarrow{timezone}Pacific\;Standard\;Time$. In KGQA, $\forall i\in[1,l],\;(e_{i-1},r_{i},e_{i})\in\mathcal{G}$. That is, it is guaranteed that all triples in gold path are contained by $\mathcal{G}$.
@@ 4.1 Task Introduction
IKGQA differs from KGQA in that, in IKGQA, $\exists i\in[1,l],\;(e_{i-1},r_{i},e_{i})\notin\mathcal{G}$. That is, it doesn’t guarantee that all triples in gold path are contained by $\mathcal{G}$. For example, the triple (Cupertino, timezone, Pacific Standard Time) in $w_{g}$ may not be contained by $\mathcal{G}$. Therefore, models need to recall their inherent knowledge or reasoning from subgraph information.
@@ 4.2 Datasets Construction
We construct two IKGQA datasets based on two widely used KGQA datasets: WebQuestionSP (WebQSP) Yih et al. (2016b) and Complex WebQuestion (CWQ) Talmor and Berant (2018). Both datasets use Freebase Bollacker et al. (2008) as their background KG. To simulate incomplete KG, we delete some crucial triples, which appear in the gold relation path, of each question from the original KG. By doing this, simple semantic parsing methods almost fail to obtain the correct answers (except for some scenarios with multiple golden paths). In order to save computational cost, we selected the first 1,000 samples of these two datasets for constructing IKGQA questions.
For simplicity, we only consider one topic entity here. The process of generating crucial triples of a question is as follows:
  1. 1.
    Extract the topic entity $e_{q}$ from the question $q$ by LLMs.
  2. 2.
    Execute the SPARQL question $s_{q}$ (given in the datasets) on the background KG $\mathcal{G}$, and get all involved triples.
  3. 3.
    Filter property node (e.g., time, text, height), then convert the involved triples into a subgraph $g$.
  4. 4.
    Utilize breadth first search (BFS) to find the shortest path $p$, from topic entity $e_{q}$ to answer entity $a_{q}$.
  5. 5.
    Sample $k$ edges in the path $p$ as the crucial triples in the question.

### v2

# Generate-on-Graph: Treat LLM as both Agent and KG for Incomplete Knowledge Graph Question Answering

@@ Abstract
To address the issues of insufficient knowledge and hallucination in Large Language Models (LLMs), numerous studies have explored integrating LLMs with Knowledge Graphs (KGs). However, these methods are typically evaluated on conventional Knowledge Graph Question Answering (KGQA) with complete KGs, where all factual triples required for each question are entirely covered by the given KG. In such cases, LLMs primarily act as an agent to find answer entities within the KG, rather than effectively integrating the internal knowledge of LLMs and external knowledge sources such as KGs. In fact, KGs are often incomplete to cover all the knowledge required to answer questions. To simulate these real-world scenarios and evaluate the ability of LLMs to integrate internal and external knowledge, we propose leveraging LLMs for QA under Incomplete Knowledge Graph (IKGQA), where the provided KG lacks some of the factual triples for each question, and construct corresponding datasets. To handle IKGQA, we propose a training-free method called Generate-on-Graph (GoG), which can generate new factual triples while exploring KGs. Specifically, GoG performs reasoning through a Thinking-Searching-Generating framework, which treats LLM as both Agent and KG in IKGQA. Experimental results on two datasets demonstrate that our GoG outperforms all previous methods.
@@ 1 Introduction
Figure 1: Comparison between three Question Answering tasks: (a) LLM only QA, (b) Knowledge Graph QA (KGQA), (c) Incomplete Knowledge Graph QA (IKGQA), where the triple (Cupertino, timezone, Pacific Standard Time) is missing. The yellow and red nodes represent topic and answer entity, respectively.
@@
Large Language Models (LLMs) Brown et al. (2020); Bang et al. (2023) have made great success in various natural language processing (NLP) tasks. Benefiting from extensive model parameters and vast amounts of pre-training corpus, LLMs can solve complex reasoning tasks through prompting engineer and in-context learning Dong et al. (2023), without fine-tuning for specific tasks.
@@
However, LLMs still suffer from insufficient knowledge and hallucination issues  Huang et al. (2023); Li et al. (2023a), as shown in Figure 1 (a). To mitigate those issues, many methods that incorporate LLM with Knowledge Graphs (KGs) Ji et al. (2021) have been proposed Pan et al. (2023), where KGs provide accurate factual knowledge in triple format, while LLMs provide strong language processing and knowledge integration ability. These works can be roughly divided into two categories, as shown in Figure 2: (1) Semantic Parsing (SP) methods Li et al. (2023c); Luo et al. (2024), which use LLMs to convert natural language questions to logical queries, and then obtain answers by executing these logical queries on KGs. (2) Retrieval Augmented (RA) methods Li et al. (2023d), which retrieve information related to the question from KGs as external knowledge to guide LLMs to generate the answers.
Figure 2: Three paradigms for combining LLMs with KGs.
@@
Semantic parsing methods exclusively treat LLMs as parser, which depend heavily on KGs’ quality and completeness Sun et al. (2023). Although retrieval augmented methods claim to solve the drawbacks of semantic parsing methods and obtain good performance on conventional Knowledge Graph Question Answering (KGQA) Yih et al. (2016a), it is still hard to verify whether they really integrate knowledge from KGs and LLMs. One crucial reason is that, in conventional KGQA tasks, the factual triples required for each question are entirely covered by the KG. For example, for the question "What is the timezone of the area where Apple headquarters is located?" in Figure 1 (b), the LLMs only need to start from "Apple headquarters", sequentially choose the relation predicates "located_in" and "timezone" to find the answer. That means, in this scenario, LLMs only need to ground the relationship mentioned in the question to the specific relation predicates in the KG to reach the answer entity "Pacific Standard Time" without really integrating internal and external knowledge.
However, on the one hand, KGs are often incomplete to cover all the knowledge required to answer questions in real-world scenarios. For example, for the same question in Figure 1 (c), the crucial triple (Cupertino, timezone, Pacific Standard Time) does not exist in the KG. On the other hand, LLMs contain rich knowledge content and possess powerful reasoning ability. For example, LLMs usually know the time zone of a city. This raises the research question: Can LLMs be combined with incomplete KGs to answer complex questions?
@@
To answer this question, in this paper, we first propose a new benchmark, which utilizes LLMs for QA under incomplete KG (IKGQA), to simulate realistic scenarios. We construct the IKGQA datasets based on existing public KGQA datasets and simulate KGs with varying degrees of incompleteness by randomly dropping triples according to different probabilities. Unlike conventional KGQA, the corresponding KG in IKGQA does not encompass all the factual triplets required for each question. This means that semantic parsing methods may fail to retrieve the final answer even generating the correct SPARQL query  [^1] Semantic parsing methods always parse ”timezone” into to ”timezone” rather than ”located_in -¿ timezone” because of the training set, more details can be found in Appendix A.. Besides, previous retrieval augmented methods also can’t perform well under incomplete KGs, as they still heavily rely on the retrieved paths, more details are in Appendix B. Compared to KGQA, IKGQA holds greater research significance for the following reasons: (1) It is closer to real-world scenarios where the given KG is incomplete to answer users’ questions. (2) It can better evaluate the ability of LLMs to integrate the internal and external knowledge.
We also propose a novel method called Generate-on-Graph (GoG) for IKGQA, as illustrated in Figure 2 (c), which not only treats LLM as an agent exploring the given KGs to retrieve relevant triples, but also as a KG to generate additional factual triples for answering this question. Specifically, GoG adopts a Thinking-Searching-Generating framework, consisting of three main steps: (1) Thinking: LLMs decompose the question and determine whether to conduct further searches or generate relevant triples based on the current state. (2) Searching: LLMs use pre-defined tools, such as a KG engineer executing SPARQL queries, to explore the KGs and filter out irrelevant triples. (3) Generating: LLMs use its internal knowledge and reasoning abilities to generate required new factual triples based on the explored subgraph and verify them. GoG will repeat these steps until obtaining adequate information to answer the question. The codes and data are available at https://github.com/YaooXu/GoG.
@@
    We propose leveraging LLMs for QA under incomplete KG (IKGQA) to better evaluate LLMs’ ability, and construct corresponding IKGQA datasets based on existing KGQA datasets.
@@
    We propose Generate-on-Graph (GoG), which uses the Thinking-Searching-Generating framework, to address IKGQA.
@@
    Experimental results on two datasets show the superiority of GoG, and demonstrate that LLMs can be combined with incomplete KGs to answer complex questions.
Figure 3: Comparison of three methods in solving IKGQA: (a) Semantic parsing based method (e.g., ChatKBQA Luo et al. (2024)), (b) Path retrieval method (e.g., ToG Sun et al. (2023)), (c) The proposed GoG with Thinking-Searching-Generating framework.
@@ 2 Related Work
Question Answering under Incomplete KG. Some previous works Saxena et al. (2020); Zan et al. (2022); Zhao et al. (2022); Guo et al. (2023) attempt to train KG embeddings to predict answers by similarity scores under incomplete KG. Compared to these previous KGE-based works, we propose leveraging LLMs for QA under incomplete KG to study whether LLMs can integrate internal and external knowledge well.
Unifying KGs and LLMs for KGQA. Various methods have been proposed to unify KGs and LLMs to solve KGQA, these methods can be classified into two categories: Semantic Parsing (SP) methods and Retrieval Augmented (RA) methods. SP methods transform the question into a structural query using LLMs. These queries can then be executed by a KG engine to derive answers based on KGs. These methods generate the drafts as preliminary logical forms first, and then bind the drafts to the executable ones with entity and relation binders, such as KB-BINDER Li et al. (2023c) and ChatKBQA Luo et al. (2024). However, the effectiveness of these methods relies heavily on the quality of the generated queries and the completeness of KGs. RA methods retrieve related information from KGs to improve the reasoning performance Li et al. (2023b). ToG Sun et al. (2023) treats the LLM as an agent to interactively explore relation paths step-by-step on KGs and perform reasoning based on the retrieved paths. RoG Luo et al. (2023) first generates relation paths as faithful plans, and then use them to retrieve valid reasoning paths from the KGs for LLMs to reason. Readi Cheng et al. (2024) generates a reasoning path and edit the path only when necessary. Salnikov et al. propose "generate-then-select" method that first uses LLMs to generate answers directly, and then constructs subgraphs and selects the subgraph most likely to contain the correct answer.
Our GoG belongs to retrieval augmented methods, we also utilize the knowledge modeling ability of LLMs, which is also similar to GAG Yu et al. (2023).
LLM reasoning with Prompting. Many works have been proposed to elicit the reasoning ability of LLMs to solve complex tasks through prompting Wei et al. (2023); Khot et al. (2023). Complex CoT Fu et al. (2023) creates and refines rationale examples with more reasoning steps to elicit better reasoning in LLMs. Self-Consistency Wang et al. (2023) fully explores various ways of reasoning to improve their performance on reasoning tasks. DecomP Khot et al. (2023) solves complex tasks by instead decomposing them into simpler sub-tasks and delegating these to sub-task specific LLMs. ReAct Yao et al. (2023) treats LLMs as agents that interact with the environment and make decisions to retrieve information from external source. GoG can be viewed as a fusion of ReAct and DecomP, thereby enabling a more comprehensive utilization of the diverse capabilities internal in LLMs for addressing complex questions.
@@ 3 Preliminary
In this section, we first introduce Knowledge Graphs (KGs). Then, we use symbols of KGs to describe relation path and Knowledge Graph Question Answering (KGQA).
Knowledge Graphs (KG) can be described as a set of inter-linked factual triples, i.e., $\mathcal{G}=\{(h,r,t)\in\mathcal{V}\times\mathcal{R}\times\mathcal{V}\}$, where $h,r\in\mathcal{V}$ denote the head and tail entity, $r\in\mathcal{R}$ represents the relation.
Knowledge Graph Question Answering (KGQA) is a reasoning task that aims to predict answer entities $e_{a}\in\mathcal{A}_{q}$ based on $\mathcal{G}$. Following previous work Sun et al. (2019), we call the entities mentioned in question $q$ as topic entities, denoted as $e_{t}\in\mathcal{T}_{q}$. Many datasets Talmor and Berant (2018); Yih et al. (2016b) give the standard SPARQL query of each question, which demonstrates a relation path from the topic entity $e_{t}$ to answer entity $e_{a}$. We call this path as gold relation path, denote it as $w_{g}=e_{q}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}...\xrightarrow{r_{l}}e_{a}$. For example, the gold relation path of the question in Figure 3 is $w_{g}=Apple\;Inc\xrightarrow{headquarter}Cupertino\xrightarrow{timezone}Pacific\;Standard\;Time$. In KGQA, $\forall i\in[1,l],\;(e_{i-1},r_{i},e_{i})\in\mathcal{G}$. That is, it is guaranteed that all triples in gold path are contained by $\mathcal{G}$.
@@ 4.1 Task Introduction
IKGQA differs from KGQA in that, in IKGQA, $\exists i\in[1,l],\;(e_{i-1},r_{i},e_{i})\notin\mathcal{G}$. That is, it doesn’t guarantee that all triples in gold path are contained by $\mathcal{G}$. For example, the triple (Cupertino, timezone, Pacific Standard Time) in $w_{g}$ may not be contained by $\mathcal{G}$. Therefore, models need to recall them from LLMs or reasoning from subgraph information.
@@ 4.2 Datasets Construction
At present, there are no IKGQA datasets readily available. In this paper, to promote relevant research, we construct two IKGQA datasets based on two widely used KGQA datasets: WebQuestionSP (WebQSP) Yih et al. (2016b) and Complex WebQuestion (CWQ) Talmor and Berant (2018). Both datasets use Freebase Bollacker et al. (2008) as their background KG. To simulate incomplete KGs, we randomly delete some crucial triples, which appear in the gold relation path, for each question from the original KG. By doing this, simple semantic parsing methods almost fail to obtain the correct answers. In order to save computational costs, we randomly select 1,000 samples of these two datasets for constructing IKGQA questions.
The process of generating crucial triples of a question is illustrated in Algorithm 1.
Input: SPARQL query $s_{q}$, KG $\mathcal{G}$, probability $p$
Output: Dropped crucial triples list $L$
1
2Initialize $L\leftarrow[]$, $filtered\_triples\leftarrow[]$;
3
[…]
5
[…]
7
8// Filter property node (e.g., height, text)
[…]
10
11for each $t$ in $filtered\_triples$ do
12       $r$leftarrow generate_random_float();
13       if $r\leq p$ then
14             $L$.add($t$)
15       end if
16
17 end for
18Return $L$;
19
Algorithm 1 Obtaining crucial triples of the question $q$